OpenStudy (anonymous):

SAT question: if a^5=9 and a^4=(4/x), which of the following is an expression for a in terms of x?

OpenStudy (anonymous):

a. 9/4x b. 36/x c. 9x-4/x d. 9x/4 e. x/36

OpenStudy (anonymous):

$$a^5=9\\a\cdot a^4=9\\a\cdot\frac4x=9\\a=\frac{9x}4$$

OpenStudy (anonymous):

@oldrin.bataku Thanks, I appreciate it a lot hah :)

OpenStudy (anonymous):

@oldrin.bataku Shouldn't it be flipped though?

OpenStudy (anonymous):

Dumb question, but yeah

OpenStudy (anonymous):

@sakigirl we flipped it from \(4/x\) to \(x/4\) when multiplying both sides, hence why we get left with \(a=9x/4\)

OpenStudy (jhannybean):

There are never any dumb questions lmao. this is on the SAT after all, right?

OpenStudy (anonymous):

@Jhannybean ahaha, thanks for making me feel better